Opened 9 years ago

Closed 3 years ago

#329 closed flight processing (fixed)

GB07/12, flight day 099/2010, Debden

Reported by: benj Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/uk/GB07_12-2010_099_Debden

Data arrived from ARSF via network transfer on 15th April 2010.

Scientific objective: Monitoring of peat fires and quantification of pyrogenic carbon conversion

Priority: a4l

PI: Martin Wooster

This is the post-burning flight relating to #323 and #324

Sensors:

  • Eagle : Delivered 13/08/2010
  • Hawk (20/12/2010)
  • Leica LIDAR : Delivered 13/08/2010
  • Digital Photography (07/09/2010)

Change History (30)

comment:1 Changed 9 years ago by benj

  • Milestone set to 2010 data processing completion

comment:2 Changed 9 years ago by benj

Eagle has an extra blank line collected near Nottingham (they were going to do another collection and were put off by the weather). I have removed line 099-10-1.* from the Eagle directory (it's static, presumably running with the lens cap on because they weren't actually trying to collect data). I have also removed SWIR099-10-12.nav from the Hawk directory - it has no associated raw (or other) file.

comment:3 Changed 9 years ago by pmlvis10

IPAS navigation for this project has been processed.

comment:4 Changed 9 years ago by ella

Looks like the applanix navigation processing has been done. Saved in arsf Samba workspace.

comment:5 Changed 9 years ago by anch

  • Owner set to anch
  • Status changed from new to accepted

Hyperspectral processed

comment:6 Changed 9 years ago by knpa

Delivery checked.

All checks out apart from:

  • 5 Readmes - delete all apart form pdf.
  • Readme contents section - flightline named in right-hand side column should be name from logsheet e.g. '020'
  • 'Quality issues...' Section - "Digital Elevation Modeland tested" - should be space between "Model" and "and"
  • Formatting on azgcoor command incorrect, put on correct lines and make sure to escape newlines with '\'
  • eagle lev1 images look strangely squashed...need to investigate this.

Have not checked hawk due to current bandsetting issue. Be sure to get this checked before delivering.

comment:7 Changed 9 years ago by benj

Timing offsets taken from config file

Line Eagle SCT Hawk SCT
3 0.94 0.00
4 1.70 0.04
5 1.20 0.00
6 1.01 0.00
7 1.00 0.53
8 0.42
9 0.01 1.09
10 0.01
11 1.17
12 0.04

comment:8 Changed 9 years ago by jb362

Beginning LiDAR processing

comment:9 Changed 9 years ago by jb362

Abandoned LiDAR processing!

comment:10 Changed 9 years ago by benj

Why abandoned?

comment:11 Changed 9 years ago by jb362

Abandoned because I am making the fixes on my delivery checked projects, then starting the hyperspectral processing in the unpacking folder so will leave this for the next person to do.

comment:12 Changed 9 years ago by anch

I have started the LIDAR processing for this project.

.LAS files have been created in ALSPP and I am now doing QC in lag

comment:13 Changed 9 years ago by anch

LIDAR processing of this project is complete and ready for checking.

I have added it to the delivery checking list.

comment:14 Changed 9 years ago by jb362

Starting LiDAR delivery check

comment:15 Changed 9 years ago by jb362

LiDAR delivery check looks fine.

comment:16 Changed 9 years ago by anch

Eagle hyperspectral and LiDAR data have been dispatched on disk.

The data was put on the same disk as project GB08/19, as they were both for the same PI.

Eagle and LiDAR have been rsynced.

comment:17 Changed 9 years ago by anch

  • Description modified (diff)

comment:18 Changed 9 years ago by emca

Camera images processed and ready for delivery check at
~airborne/workspace/GB07_12-2010_099_Debden/delivery/20100820/GB07_12

Removed images 1-9 as they did not cover the study area

comment:19 Changed 9 years ago by jb362

Delivery checking camera.

comment:20 Changed 9 years ago by jb362

Camera delivery looks fine.

comment:21 Changed 9 years ago by mark1

  • Description modified (diff)

Photography data dispatched to PI on 07/09/2010

comment:22 Changed 9 years ago by mark1

Hawk has been delivery checked. Read me needs regenerating to remove files listed that are not present. Otherwise OK to deliver.

comment:23 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data sent 08/12/2010

comment:24 Changed 9 years ago by anch

Hawk data reprocessed and added to delivery 20101207

comment:25 Changed 9 years ago by mark1

Hawk delivery checked and looks OK. Ready to re-deliver.

comment:26 Changed 9 years ago by emca

Hawk data re-delivered 20/12/2010

comment:27 Changed 9 years ago by emca

  • Description modified (diff)

comment:28 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from anch to knpa
  • Status changed from accepted to assigned
  • Summary changed from GB07/12, flight day 99/2010, Debden to GB07/12, flight day 099/2010, Debden

Beginning archiving.

comment:29 Changed 3 years ago by dac

Archiving

Only reprocessed hyperspectral data available from NEODC. Started uploading LiDAR, camera and raw data.

comment:30 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Processed and raw data available from NEODC (processsed: http://browse.ceda.ac.uk/browse/neodc/arsf/2010/GB07_12/GB07_12-2010_099_Debden).

Marking as closed.

Note: See TracTickets for help on using tickets.